OpenStudy (anonymous):

step by step please 4(x+5) = 3(x-4) +x

OpenStudy (anonymous):

4x+20 = 3x-12+x 4x+20 = 4x-12 20=-12 that's not true so no solutions

OpenStudy (anonymous):

4x+20 = 3x-12+x 4x-3x-x = -12-20 0 = -32